OpenStudy (anonymous):

Find the area of a ring bounded by circles with radii 11 cm and 8 cm. (Show the solution. Thanks)

OpenStudy (campbell_st):

 its the area of the big circle - area of the small circle \[A = \pi \times 11^2 - \pi \times 8^2 \]

OpenStudy (anonymous):

That's the answer? :">

OpenStudy (campbell_st):

that will give you the answer... you need to evaluate it

OpenStudy (anonymous):

Solve it more..

OpenStudy (anonymous):

What is 11*11 = ??

OpenStudy (campbell_st):

here is a calculator for you http://web2.0calc.com/

OpenStudy (anonymous):